# Rate of epidemic spreading on complex networks

### v1

@@ Abstract
The initial phase of an epidemic is often characterised by an exponential increase in the number of infected individuals. In well-mixed populations, this exponential increase is controlled by the basic reproduction number $R_{0}$ and the distribution of times between consecutive infection along an infection chain. However, we are still lacking a general understanding of how epidemics spread when individual interactions form a complex network. Here, we derive an expression for the rate of exponential spread of an epidemic on a complex network. We find that this rate is affected by the degree distribution, the network assortativity, and the level of clustering. Our result holds for a broad range of networks, aside from networks with very broad degree distribution, where no clear exponential regime is present. The theory presented in this paper bridges the gap between classic epidemiology and the theory of complex network, with broad implications for model inference and policy making.
@@
Epidemic modelling permits to rationalise evidences, test hypotheses, and inform disease control and other policy making. Epidemiological models should incorporate the key features that determine the epidemic spreading. One such feature is the fact that individuals are arranged in social structures [1, 2]. These social structures play a crucial role in the spreading of an epidemic, and in the effectiveness of intervention strategies [3, 4]. In these social structures, the number of contacts per individual is highly variable. For example, at the early stages of COVID-19, the individuals with most contacts contributed disproportionately to the epidemic [5, 6].
@@
A convenient way of modelling social structures is by means of complex networks [7], in which nodes represent individuals and links signify that pairs of individuals are in close contact. A success of network theory has been to determine the epidemic threshold on networks [8], i.e., the transition point from a localised outbreak to a widespread epidemic. This theory has shown that networks whose degree distribution has a diverging variance due to heavy tails [9, 10] are extremely vulnerable to epidemics. A paradigmatic example is found in the Barabási-Albert model [11], where a network is constructed by preferentially attaching new nodes to existing nodes with high degree. Another property of complex networks which is relevant for epidemic spreading is clustering, which is a measure of how commonly nodes tend to form tightly connected communities. Many networks in nature possess both fat tails in their degree distribution and high level of clustering [12].
@@
A second key feature is the fact that the infectiousness of an individual, i.e., the rate at which an infected individuals infects its contacts, tends to change in time in a way that is particular to each disease. The underlying reason is that the infectiousness incorporates multiple evolving factors, such as the viral load of an individual, its social behaviour, and possibly environmental factors [13]. This time-dependence profoundly impacts the way a disease spreads.
@@
Social structures and individual infectiousness are the key factors in determining the initial stage of an epidemic, in which most individuals are susceptible and the number of infected individuals usually grows exponentially. This stage is of central importance for novel epidemics, or during emergence of novel variants, often requiring rapid policy making. At later stages, other factors become relevant, such as individual recovery time, their behavioural change as a consequence of the epidemic [14], and the possibility of being reinfected.
@@
Despite the vast body of theory on epidemics in networks [10], results on the initial exponential regime are surprisingly sparse. A recent theory estimates the average time it takes for an epidemic to reach an individual at a certain distance from patient zero [15]. The spread of an epidemic on a network can be exactly solved [16], but this calculation becomes practically unfeasible unless the network is very small. Other approaches avoid a detailed description of social structures, and effectively model the heterogeneity of the population by assigning different risk propensities to individuals [17, 18].
In this paper, we directly estimate the exponential rate of epidemic spreading on a complex network. Our approach is based on classic ideas of theoretical epidemiology, that we extend to complex networks.
@@
where the basic reproduction number $R_{0}$ represents the average number of secondary infections by each infected individual, and the dot denotes a time derivative. The function $w(\tau)$ is the infectiousness, defined as the probability density that a given secondary infection occurs after a time $\tau$ from the primary one. We assume to be in the exponential regime, $I(t)\sim e^{\Lambda t}$, where $\Lambda$ is the rate of exponential spread. Substituting this assumption into Eq. (1) we obtain the Euler-Lotka equation
@@
Figure 1: Epidemic spreading on complex networks. a: Epidemic on a network. Nodes can be infected (red) or healthy (blue). Infected individuals spread the infection at a rate $\lambda(\tau)$ to each of their neighbours, to whom they are connected via links. Individuals may have different numbers of neighbours. Due to the presence of cycles, some neighbours may have already been infected by someone else, so that the transmission does not result in an infection. b: Infected individuals spread the disease at a time-dependent rate $\lambda(\tau)$, where $\tau$ is the time since they were infected. The resulting infection times are distributed according to $\psi(\tau)$. c: Example of an epidemic trajectory starting with one infected individual. After an outbreak, the epidemic spreads at an exponential spreading rate $\Lambda$. The exponential phase ends in a saturation phase, where a sizeable fraction of the individuals are already infected. d: Epidemic tree. The epidemic starts with a first infected individual (patient zero) and resulting infections are represented by continuous lines. Dashed lines represents different generations.
@@
We define the basic reproduction number $R_{0}$ for a network as the average number of contacts of an infected individual in the early stage of the epidemic, excluding the one from which the infection came from. We also define the effective reproduction number $R$ as the average number of infections caused by an infected individual. In contrast with $R_{0}$, $R$ does not include contacts that are infected by other individuals. This difference is crucial in clustered networks.
@@
On the one hand, assuming that $R$ remains constant between generations, the average number of infected individuals at the $n$th generation is equal to $Z_{n}\sim R^{n}$, up to the leading order. On the other hand, the number of infected individuals $Z_{n}$ scales as the $n$-th power of the reproduction matrix (see Methods). This permits us to identify $R$ as the leading eigenvalue of the reproduction matrix. The associated right eigenvector $v_{k^{\prime}}$ can be interpreted as the degree distribution of the infected individuals.
For large networks, diagonalising the reproduction matrix can be unwieldy. In the simple case where there is no clustering $(m_{kk^{\prime}}=0$ for all links) and the degrees of connected nodes are uncorrelated, we have
@@
where the bar represents an average over the degree distribution. Equation (4) is a classic result for epidemic spreading on networks [9, 21, 8]. We seek to extend it to the general case of clustered and correlated networks. We quantify degree correlations by the assortativity coefficient $r$. When $r>0$, nodes with high degree tend to connect together whereas when $r<0$, they tend to connect to lower degree nodes. We find that, for small clustering coefficients and degree correlation, the basic reproduction number $R$ is approximated by
@@
where $c_{k}$ is the probability that two neighbours of a node of degree $k$ are themselves connected, $m_{1}=\overline{k(k-1)c_{k}}/\overline{k}$ is the average number of triangles a link is part of, and $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$ (see Methods). The reason for calling $R_{\mathrm{pert}}$ the expression in Eq. (5) is that this is obtained by a perturbative expansion in $r$ (see Methods).
We compare the expression for $R_{\mathrm{pert}}$ given in Eq. (5) to the maximum eigenvalue $R$ of the matrix $M_{kk^{\prime}}$ for different network models, see Table 1. We find that the relative difference between $R$ and $R_{\mathrm{pert}}$ is within $2\%$ for all the networks we considered, except for Barabási-Albert with $m=1$ ($20.6\%$) and $m=2$ ($7.1\%$). We shall scrutinize the reason for this discrepancy later in the paper. Here and in the following, when computing $R_{\mathrm{pert}}$ using Eq. (5) for a given network, we always interpret the averages as empirical averages. Because of this choice, moments of the degree distribution are always finite, even in very heterogeneous networks. In practical cases, Eq. (5) predicts that the reproduction number $R_{\mathrm{pert}}$ increases with the assortativity $r$. Accordingly, assortativity lowers the epidemic threshold, as previously known for the unclustered case [22]. However, the coefficient multiplying $r$ in Eq. (5) is usually rather small, except when the degree distribution has heavy tails. In contrast, clustering systematically lowers $R_{\mathrm{pert}}$, regardless of the degree distribution. Therefore, approximating the reproduction number with $R_{0}$ leads to an overestimate in the presence of clustering.
@@
Network Assortativity $r$ Clustering $\overline{c}$ $R_{0}$ $R_{\mathrm{pert}}^{(r=0)}$ $R_{\mathrm{pert}}^{(\overline{c}=0)}$ $R_{\mathrm{pert}}$ $R$
$\text{BA}_{1}$0.011 0 13.9 13.9 6.6 6.8 8.5
$\text{BA}_{2}$0.008 0.0001 20.4 20.3 14.2 14.4 15.5
$\text{BA}_{12}$0.004 0.0003 85.4 83.6 80.4 79.5 80.8
ER6 0.001 0.000 6.0 6.0 6.0 6.0 6.0
[…]
WS8,2 -0.022 0.334 7.2 4.9 7.0 4.9 4.9
WS6,2 -0.029 0.315 5.2 3.6 5.0 3.6 3.6
$\text{LN}_{4,1}\;\alpha=0.5$ 0.007 0.161 3.7 3.2 3.8 3.2 3.2
Parameters associated with the epidemic spreading for different models of complex networks: BAm: Barabási-Albert model with parameter $m$. $\text{ER}_{d}$: Erdős-Rényi with mean degree $d$. WSk,p: Watts-Strogatz starting from a 1D lattice with $k$ nearest neighbours and rewiring probability $p$. $\text{LN}_{4,1}\;\alpha=0.5$ is generated from a modified version of the configuration model [23] with a Lognormal degree distribution with mean $4$ and variance $1$. This modified algorithm allows us for tuning the clustering coefficients as $c_{k}\propto(k-1)^{-\alpha}$ for a chosen $\alpha>0$. $\text{ER}_{10}\;\alpha=0.2$: A Erdős-Rényi network in which we imposed clustering using the same algorithm. We have defined the average clustering coefficient $\bar{c}=\sum_{k}c_{k}P(k)$. The reproduction number $R$ is the numerically computed leading eigenvalue of Eq. (3), whereas we obtain $R_{\mathrm{pert}}$ by evaluating Eq. (5). To assess the impact of degree correlations and clustering, we also compute $R_{\mathrm{pert}}$ for $\overline{c}=0$.
@@
Equation (6) extends Eq. (1) to complex networks. It is linear because of the assumption that the epidemic is in the initial stage and thus far from saturation. We assume exponential growth, $\dot{I}_{k^{\prime}}(u)\propto v_{k^{\prime}}e^{\Lambda t}$. Substituting this assumption into Eq. (6) we obtain
$\frac{1}{R}=\left\langle e^{-\Lambda\tau}\right\rangle\,.$ (7)
We call Eq. (7) the network Euler-Lotka equation (see section II in SI for an alternative derivation based on large deviation theory). Although the course of infectiousness and the network topology conspire in determining the spreading rate $\Lambda$, these two aspects appear as separated in Eq. (7). Indeed, the left hand side of Eq. (7) is uniquely determined by the network topology, while the right hand side only depends on $\psi(\tau)$ at given $\Lambda$. Equation (7) predicts that the epidemic threshold, i.e the condition $\Lambda=0$, is obtained by setting $R=1$, independently of $\psi(\tau)$.
We shall use Eq. (7) as an implicit relation for the unknown $\Lambda$. Using the first two moments of the generation-time distribution, we also obtain an explicit, although approximated expression:
@@
where $\sigma^{2}=\langle\tau^{2}\rangle-\langle\tau\rangle^{2}$, see section II-B in SI. An alternative is to combine Eq. (7) with the Jensen inequality, that states that $\langle e^{-\Lambda\tau}\rangle\geq e^{-\Lambda\langle\tau\rangle}$. This leads to a lower bound on the exponential spreading rate of the epidemic:
@@
The network Euler-Lotka equation successfully predicts the rate of epidemic spreading on synthetic complex networks, Fig. 2-a and Fig. 2-b. Our battery of tests includes a broad range of models characterised by different degree distribution, assortativity, and clustering. For each model, we ran simulations using three different infectivity functions, see Fig. 2-c. The choice of the infectivity function qualitatively affect the behaviour of the epidemic. In particular, for peaked distributions, the exponential growth appears modulated by oscillations, since early generations of the epidemic are nearly synchronised. To extrapolate the leading exponential behaviour from these curves, we fit them to a generalised logistic function, which appear to capture well finite-size effects (see Methods). These fits also permits to define a finite-size parameter $\eta\in[0,1]$, which quantifies the relative impact of finite-size effects on the exponential spreading rate at short times (see Methods and section III in SI). The difference between the exponential spreading rate estimated from the simulations and from the network Euler-Lotka equation (7) is within $6\%$ error for all networks, see Fig. 2-f, except for the Barabási-Albert network with parameter $m=1$.
@@
Figure 2: The network Euler-Lotka equation predicts the rate of epidemic spreading on synthetic networks.: Panel a: Average epidemic on a Erdős-Rényi network of size $N=10^{6}$ with $\overline{k}=4$ with Gamma distributed infection times with mean $7$ and variance $1$. The red curve is an average over 500 realisations of the epidemic. The exponential spreading rate (yellow dashed line) is obtained by fitting a generalised logistic function on the average trajectory (red dots). The black dashed line is the solution of the network Euler-Lotka equation with $R=R_{\mathrm{pert}}$ as given by Eq. (5). Panel b: Same as a, but with a Weibull distribution with shape $2$ and scale $2$ on a Watts-Strogatz network of size $N=10^{6}$. Panel c: Solution of the network Euler-Lotka equation compared with simulations for various networks. The procedure for a - b is repeated for various networks of sizes $10^{5}$ and $10^{6}$: Barabási-Albert (circles); Watts-Strogatz (diamonds); Erdős-Rényi (triangles); networks from the configuration model (CM: squares) with Lognormal and Borel degree distributions; networks from a generalisation of the configuration model (CCM: pentagons) whose assortativity and clustering coefficient have been modified (see details in section IV in SI). These simulations are repeated for three distributions: Weibull with parameter $2$ and scale $2$ (dashed green lines), Lognormal with mean $5$ and variance $3$ (solid blue lines), Gamma with mean $7$ and variance $1$ (dotted yellow lines). The black arrows point to the case of Barabási-Albert network for $m=1$. For these points, the values of the finite-size parameter are $\eta=0.68$ (Weibull), $\eta=0.245$ (Gamma), $\eta=0.400$ (Lognormal). In comparison, for $m=2$ we have $\eta=0.018$ (Weibull), $\eta=0.001$ (Gamma), $\eta=0.003$ (Lognormal). For networks other than Barabási-Albert networks, we find that on average $\eta=4.2\times 10^{-6}$ with worst case $\eta=1.2\times 10^{-5}$ (Lognormal network, Weibull distribution).
We now discuss the Barabási-Albert case in more detail. The parameter $m$ in the Barabási-Albert model represents the number of nodes a new nodes attaches to. In the case $m>1$, the network Euler-Lotka equation well predicts the spreading rate (Fig. 2f), despite these networks being scale-free, disassortative, and including loops. In contrast, the error in the estimate of the spreading rate for $m=1$ is approximately $15\%$. There are two possible causes for this fact. The first is that the exponential regime for $m=1$ is not clearly defined, as revealed by the finite-size parameter (see Fig. 2). The second is that, for $m=1$, the validity of the first-order perturbative expansion in $r$ might break down (see section V-A in SI). Indeed, a similar problem occurs when computing the average path length of the Barabási-Albert model for small $m$ [24]. This would also explain the large difference between $R_{\mathrm{pert}}$ and the leading eigenvalue of the reproductive matrix (see Table 1).
@@
Figure 3: Epidemic spreading rate on real networks. We simulate epidemic spreading on real-world networks from ICON [25], SNAP [26] and KONECT [27]. All considered networks are undirected and without multiple links or self-connections. Their size varies from $N=10^{3}$ to $N=10^{6}$ (see section IV in SI). We simulate $500$ epidemics on each of these empirical networks with lognormally distributed infection times with mean $5$ and variance $3$. Panels a and b show the average trajectory for two different networks along with a representation of the actual network. Similarly to Figure 2, for each network, the exponential spreading rate (blue dashed-dotted line) is obtained by fitting a generalised logistic function to the average trajectory (red dots). The black dashed line represents the slope predicted by the network Euler-Lotka equation. Panel c: Exponential spreading rate for all empirical networks. The solid line is the solution of the Euler-Lotka equation, and the markers are the numerical fits. The colour of the marker denotes the value of the finite-size parameter $\eta$.
Empirical networks
@@
We apply the Euler-Lotka equation to a large set of social, biological, technological, and transportation networks, see Fig. 3. These real-world networks exhibit more complex structures than network models, often characterised by tightly linked communities, strong degree-correlation and broad tails in the degree distribution. Therefore, they constitute a more challenging test. Surprisingly, the network Euler-Lotka equation holds well, see Fig. 3-a and Fig. 3-b for examples.
@@
In general, we find that the prediction of the network Euler-Lotka equation well reproduces the fitted spreading rate whenever the finite-size parameter $\eta$ detects the presence of a clear exponential regime see Fig. 3-c. Despite some of these networks being highly clustered, using $R_{\mathrm{pert}}$ as expressed in Eq. (5) does not perform worse than $R$ (See tables section IV-E in SI). In Fig. 3-c, we have omitted three road networks (California, Texas, and Pennsylvania). For these networks, the generalized logistic fit fails to converge. However, a function of the form $\dot{I}(t)\sim t^{\beta}$ well fits the epidemic trajectories on these networks (see section IV-D in SI). Likely, the fact that these networks are embedded in a two-dimensional physical space prevents an exponential epidemic spread.
To better understand the cases where the error in the predicted spreading rate is large, we first scrutinise whether these networks are heavy-tailed. We quantify this by fitting the tail of the degree distribution by a power law of the form $k^{-\gamma}$ multiplied by a slowly varying function, see [28] and section VI in SI. We find large errors only when the finite-size parameter is large and $\gamma\leq 3$, see Fig. 4-a. In contrast, the error is weakly affected by the strength of clustering, see Fig. 4-b.

### v2

@@ Abstract
The initial phase of an epidemic is often characterised by an exponential increase in the number of infected individuals. In a well-mixed population, this exponential increase is controlled by the basic reproduction number $R_{0}$ and the distribution of times between consecutive infection along an infection chain. However, real populations are characterised by a complex network of interactions, and are thus far from being well mixed. Here, we derive an expression for the rate of exponential spread of an epidemic on a complex network. We find that this rate is affected by the degree distribution, the network assortativity, and the level of clustering. Our result holds for a broad range of networks, apart from networks with very broad degree distribution, where no clear exponential regime is present. The theory presented in this paper bridges the gap between classic epidemiology and the theory of complex network, with broad implications for model inference and policy making.
@@
Epidemic modelling permits rationalisation of evidences and can provide crucial information for policy making. Epidemiological models should incorporate the key features that determine epidemic spreading. One such feature is the arrangement of individuals in social structures [1]. These social structures play a crucial role in the spreading of an epidemic, and in the effectiveness of intervention strategies [2, 3]. In these social structures, the number of contacts per individual is highly variable. For example, at the early stages of COVID-19, the individuals with most contacts contributed disproportionately to the epidemic [4, 5].
@@
A convenient way of modelling social structures is by means of complex networks [6], in which nodes represent individuals and links signify pairs of individuals that are in close contact. A success of network theory has been to determine the epidemic threshold on networks [7, 8, 9], i.e., the transition point from a localised outbreak to a widespread epidemic. This theory has shown that networks whose degree distribution has a diverging variance due to heavy tails [7, 10, 9] are extremely vulnerable to epidemics. A paradigmatic example is found in the Barabási-Albert model [11], where a network is constructed by preferentially attaching new nodes to existing nodes with high degree. Another property of complex networks which is relevant for epidemic spreading is clustering, which is a measure of how commonly nodes tend to form tightly connected communities. Many networks in nature possess both fat tails in their degree distribution and high level of clustering [12].
@@
Another key feature of an epidemic is its infectiousness, i.e., the rate at which an infected individual infects its contacts. The infectiousness incorporates multiple factors, such as the viral load of an individual, its social behaviour, and possibly environmental factors [13]. Because of changes in these factors, the infectiousness of an individual tends to change in time in a way that is particular to each disease. This time-dependence profoundly impacts the way a disease spreads.
@@
Social structures and individual infectiousness are the key factors in determining the initial stage of an epidemic, in which most individuals are susceptible and the number of infected individuals usually grows exponentially. This stage is of central importance for novel epidemics, or during emergence of novel variants, and it often requires rapid policy making. At later stages, other factors become relevant, such as individual recovery time, their behavioural change as a consequence of the epidemic [14], and the possible occurrence of reinfections.
@@
Despite the vast body of theory on epidemics in networks [9], general results on the speed of epidemic spreading are surprisingly sparse. The rate of epidemic spreading can be calculated for unclustered, uncorrelated networks, and with constant infectiousness [15]. A recent theory estimates the average time it takes for an epidemic to reach an individual at a certain distance from patient zero [16]. The spread of an epidemic on a network can be exactly solved [17], but this calculation becomes practically unfeasible unless the network is very small. Other approaches avoid a detailed description of social structures, and effectively model the heterogeneity of the population by assigning different risk propensities to individuals [18, 19].
In this paper, we present a general theory that predicts the exponential rate of epidemic spreading on a complex network. In particular, we quantify the impact of key properties of the network, such as assortativity and clustering, on the reproduction number of the epidemic. We further link the reproduction number and the individual infectiousness with the epidemic spreading rate. We shall see that this approach leads to accurate predictions for a broad range of networks.
@@
where the basic reproduction number $R_{0}$ represents the average number of secondary infections by each infected individual, and the dot denotes a time derivative. The function $w(\tau)$ is the infectiousness, defined as the probability density that a given secondary infection occurs after a time $\tau$ from the primary one. We assume to be in the exponential regime, $I(t)\sim e^{\Lambda t}$, where $\Lambda$ is the rate of exponential spread. Substituting this assumption into Eq. (1) and taking the limit of large time, we obtain the Euler-Lotka equation
@@
Figure 1: Epidemic spreading on complex networks. a: Epidemic on a network. Individuals (nodes) can be infected (red) or healthy (blue). Infected individuals spread the infection at a rate $\lambda(\tau)$ to each of their neighbours, to whom they are connected via links. Individuals may have different numbers of neighbours. Due to the presence of cycles, some neighbours may have already been infected by someone else, so that the transmission does not result in an infection. b: Infected individuals spread the disease at a time-dependent rate $\lambda(\tau)$, where $\tau$ is the time since they were infected. The resulting infection times are distributed according to $\psi(\tau)$. c: Example of an epidemic trajectory starting with one infected individual. After an outbreak, the epidemic spreads at an exponential spreading rate $\Lambda$. The exponential phase ends in a saturation phase, where a sizeable fraction of the individuals are already infected. d: Epidemic tree. The epidemic starts with a first infected individual (patient zero) and resulting infections are represented by continuous lines. Dashed lines represents different generations.
@@
The basic reproduction number $R_{0}$ for a network is the average number of links of an infected individual in the early stage of the epidemic, excluding the one from which the infection came from. We also define the effective reproduction number $R$ as the average number of infections caused by an infected individual. In contrast with $R_{0}$, $R$ does not include contacts that are infected by other individuals. This difference is crucial in clustered networks.
@@
On the one hand, assuming that $R$ remains constant between generations, the average number of infected individuals at the $n$th generation is equal to $Z_{n}\sim R^{n}$, up to the leading order. On the other hand, the number of infected individuals $Z_{n}$ scales as the $n$-th power of the reproduction matrix (see Methods). This permits us to identify $R$ as the leading eigenvalue of the reproduction matrix. The associated right eigenvector $v_{k^{\prime}}$ can be interpreted as the degree distribution of the infected individuals (see Methods).
For large networks, diagonalising the reproduction matrix can be unwieldy. In the simple case of an unclustered network $(m_{kk^{\prime}}=0$ for all links) in which the degrees of connected nodes are uncorrelated, we have
@@
where the bar represents an average over the degree distribution. Equation (4) is a classic result for epidemic spreading on networks [7, 22, 8]. We seek to extend it to the general case of clustered and correlated networks. We quantify degree correlations by the assortativity coefficient $r$ (see Methods). When $r>0$, nodes with high degree tend to connect together whereas when $r<0$, they tend to connect to lower degree nodes. We find that, for small clustering coefficients and degree correlation, the effective reproduction number $R$ is expressed by
@@
where $c_{k}$ is the probability that two neighbours of a node of degree $k$ are themselves connected, $m_{1}=\overline{k(k-1)c_{k}}/\overline{k}$ is the average number of triangles a link is part of, and $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$. Equation (5) is obtained by a perturbative expansion in $r$ (see Methods).
We compare the expression for $R_{\mathrm{pert}}$ given in Eq. (5) to the maximum eigenvalue $R$ of the matrix $M_{kk^{\prime}}$ for different network models, see Table 1. We find that the relative difference between $R$ and $R_{\mathrm{pert}}$ is within $2\%$ for all the networks we considered, except for Barabási-Albert with $m=1$ ($20.6\%$) and $m=2$ ($7.1\%$). We shall analyse the reason for this discrepancy later in the paper. Approximating $R$ with $R_{0}$ given in Eq. (4) leads to substantially larger errors, supporting that clustering and assortativity need to be taken into account to appropriately estimate $R$. Here and in the following, when computing $R_{\mathrm{pert}}$ using Eq. (5) and $R_{0}$ using Eq. (4) for a given network, we always interpret the averages as empirical averages. With this choice, moments of the degree distribution are always finite, even in very heterogeneous networks. In practical cases, Eq. (5) predicts that the reproduction number $R_{\mathrm{pert}}$ increases with the assortativity $r$. Accordingly, assortativity lowers the epidemic threshold, as previously known for the unclustered case [23]. However, the coefficient multiplying $r$ in Eq. (5) is usually rather small, except when the degree distribution has heavy tails. In contrast, clustering systematically reduces $R_{\mathrm{pert}}$, regardless of the degree distribution. For this reason, we find that $R_{0}$ overestimates $R$ in all the synthetic networks we considered.
@@
Network $R_{0}$ $R$ $R_{\mathrm{pert}}$ $R_{\mathrm{pert}}^{(r=0)}$ $R_{\mathrm{pert}}^{(\overline{c}=0)}$ Assortativity $r$ Clustering $\overline{c}$
$\text{BA}_{1}$ 13.9 8.5 6.8 13.9 6.6 -0.011 0
$\text{BA}_{2}$ 20.4 15.5 14.4 20.3 14.2 -0.008 0.0001
$\text{BA}_{12}$ 85.4 80.8 79.5 83.6 80.4 -0.004 0.0003
$\text{ER}_{6}$ 6.0 6.0 6.0 6.0 6.0 0.001 0.000
$\text{ER}_{10}\;\alpha=0.2$ 10.0 9.0 9.1 9.0 10.7 0.062 0.108
$\text{WS}_{8,2}$ 7.2 4.9 4.9 4.9 7.0 -0.022 0.334
$\text{WS}_{6,2}$ 5.2 3.6 3.6 3.6 5.0 -0.029 0.315
$\text{LN}_{4,1}\;\alpha=0.5$ 3.7 3.2 3.2 3.2 3.8 0.007 0.161
Parameters associated with the epidemic spreading for different models of complex networks. The basic reproduction number $R_{0}$ is given by Eq. (4). The effective reproduction number $R$ is the numerically computed leading eigenvalue of Eq. (3). We compare it with the analytical expression of $R_{\mathrm{pert}}$ given in Eq. (5). To assess the impact of degree correlations and clustering, we also compute $R_{\mathrm{pert}}$ for $r=0$ and for $\overline{c}=0$, where we have defined $\bar{c}=\sum_{k}c_{k}P(k)$. BAm: Barabási-Albert model with parameter $m$. $\text{ER}_{d}$: Erdős-Rényi with mean degree $d$. WSk,p: Watts-Strogatz starting from a 1D lattice with $k$ nearest neighbours and rewiring probability $p$. $\text{LN}_{4,1}\;\alpha=0.5$ is generated from a modified version of the configuration model [24] with a Lognormal degree distribution with mean $4$ and variance $1$. This modified algorithm allows us for tuning the clustering coefficients as $c_{k}\propto(k-1)^{-\alpha}$ for a chosen $\alpha>0$. $\text{ER}_{10}\;\alpha=0.2$: A Erdős-Rényi network in which we imposed clustering using the same algorithm.
@@
Equation (6) extends Eq. (1) to complex networks. It is linear because of the assumption that the epidemic is in the initial stage and thus far from saturation. We assume exponential growth, $\dot{I}_{k^{\prime}}(u)\propto v_{k^{\prime}}e^{\Lambda t}$. Substituting this assumption into Eq. (6) we obtain, for large time,
$\frac{1}{R}=\left\langle e^{-\Lambda\tau}\right\rangle\,,$ (7)
[…]
In general, we shall use Eq. (7) as an implicit relation for the unknown $\Lambda$. It is possible to explicitly express $\Lambda$ from Eq. (7), but only in an approximate way. In particular, using the first two moments of the generation-time distribution, we obtain
@@
where $\sigma^{2}=\langle\tau^{2}\rangle-\langle\tau\rangle^{2}$, see Section LABEL:SI-gaussian in SI. An alternative is to combine Eq. (7) with the Jensen inequality, that states that $\langle e^{-\Lambda\tau}\rangle\geq e^{-\Lambda\langle\tau\rangle}$. This leads to a lower bound on the exponential spreading rate of the epidemic:
@@
The network Euler-Lotka equation successfully predicts the rate of epidemic spreading on synthetic complex networks, Fig. 2-a and Fig. 2-b. Our battery of tests includes a broad range of models characterised by different degree distribution, assortativity, and clustering. For each model, we ran simulations using three different infectivity functions, see Fig. 2-c. The choice of the infectivity function qualitatively affect the behaviour of the epidemic. In particular, for peaked distributions, the exponential growth appears modulated by oscillations (Fig. 2-a), since early generations of the epidemic are nearly synchronised. To extrapolate the leading exponential behaviour from these curves, we fit them to a generalised logistic function, which appear to well capture finite-size effects (see Methods). The fits also permits to define a finite-size parameter $\eta\in[0,1]$, which quantifies the relative impact of finite-size effects on the exponential spreading rate at short times (see Methods and Section LABEL:SI-sec:eta in SI). When predicting the exponential growth rate using the network Euler-Lotka equation, we use $R_{\text{pert}}$ as defined in Eq. (5) as the reproduction number. As expected, the predictions are considerably worse if we instead use $R_{0}$, see Section LABEL:SI-sec:R_vs_R0 in SI. The difference between the exponential spreading rate estimated from the simulations and from the network Euler-Lotka equation (7) is within $6\%$ for all networks, see Fig. 2-c. The only exception is the Barabási-Albert network with parameter $m=1$ (average error $15\%$), where $m$ represents the number of nodes a new node attaches to.
@@
Figure 2: The network Euler-Lotka equation predicts the rate of epidemic spreading on synthetic networks.: Panel a: Average epidemic on a Erdős-Rényi network of size $N=10^{6}$ with $\overline{k}=4$ with Gamma distributed infection times. The red curve is an average over 500 realisations of the epidemic. We estimate the exponential spreading rate (yellow dashed line) by fitting a generalised logistic function on the average trajectory (red dots), see Methods. The black dashed line is the solution of the network Euler-Lotka Equation (7) with $R=R_{\mathrm{pert}}$ as given by Eq. (5). Panel b: Same as a, but with a Weibull distribution on a Watts-Strogatz network of size $N=10^{6}$. Panel c: Solution of the network Euler-Lotka equation compared with simulations for various networks. The procedure for a - b is repeated for various networks of sizes $10^{5}$ and $10^{6}$: Barabási-Albert (circles); Watts-Strogatz (diamonds); Erdős-Rényi (triangles); networks from the configuration model (CM: squares) with Lognormal and Borel degree distributions; networks from a generalisation of the configuration model (CCM: pentagons) whose assortativity and clustering coefficient have been modified (see Section LABEL:SI-sec:simulations in SI). These simulations are repeated for three generation time distributions: Weibull (dashed green lines), Lognormal (solid blue lines), Gamma (dotted yellow lines). The expressions of these distributions are given in SI Section LABEL:SI-sec:simulations.
There are two causes for this exception. The first is that the exponential regime for this model is not clearly defined, as revealed by the finite-size parameter (see Fig. 2). This issue is likely related with the fact that the second moment $\bar{k^{2}}$ of the degree distribution diverges in the infinite size limit. The second cause is that, for $m=1$, the first-order perturbative expansion in $r$ loses accuracy, as revealed by the difference between $R_{\mathrm{pert}}$ and the leading eigenvalue of the reproductive matrix (see Table 1).
@@
Figure 3: The network Euler-Lotka equation predicts the epidemic spreading rate on real-world networks. We simulate epidemic spreading on real-world networks from ICON [25], SNAP [26] and KONECT [27]. All considered networks are undirected and without multiple links or self-connections. Their size varies from $N=10^{3}$ to $N=10^{6}$ (see Section LABEL:SI-sec:real_networks in SI). We simulate $500$ epidemics on each of these real-world networks with lognormally distributed infection times, see SI Section LABEL:SI-sec:simulations. Panels a and b show the average trajectory for two different networks along with a representation of the actual network. Similarly to Figure 2, for each network, we obtain the exponential spreading rate (blue dashed-dotted line) by fitting a generalised logistic function to the average trajectory (red dots). The black dashed line represents the slope predicted by the network Euler-Lotka equation. Panel c: Exponential spreading rate for all real-world networks. The solid line is the solution of the Euler-Lotka equation. The markers are the numerical fits. The colour of the marker denotes the value of the finite-size parameter $\eta$.
Real-world networks
@@
We apply the network Euler-Lotka equation to a large set of social, biological, technological, and transportation networks, see Fig. 3. These real-world networks exhibit more complex structures than synthetic networks. They are often characterised by tightly linked communities, strong degree-correlation and broad tails in the degree distribution. Therefore, they constitute a more challenging test. Nevertheless, the network Euler-Lotka equation holds well, see Fig. 3-a and Fig. 3-b for examples.
@@
More generally, we find that the prediction of the network Euler-Lotka equation well reproduces the fitted spreading rate whenever the finite-size parameter $\eta$ detects a clear exponential regime, see Fig. 3-c. Although some of these networks are highly clustered, using $R_{\mathrm{pert}}$ as expressed by Eq. (5) does not perform worse than $R$ (See Section LABEL:SI-sec:R_vs_R0 in SI). In Fig. 3-c, we have omitted three road networks (California, Texas, and Pennsylvania). For these networks, the generalised logistic fit fails to converge. However, a function of the form $\dot{I}(t)\sim t^{\beta}$ well fits the epidemic trajectories on these networks (see Section LABEL:SI-power-law-fit in SI). Likely, the fact that these networks are embedded in a two-dimensional physical space prevents an exponential epidemic spread.
For networks in which our fit detects small finite-size effects ($\eta<0.5$), the predictions of the network Euler-Lotka equation using $R_{\text{pert}}$ are more accurate than those using $R_{0}$. However, for larger values of $\eta$, $R_{0}$ performs slightly better on average (see Section LABEL:SI-sec:R_vs_R0 in SI). Even in these cases, $R_{\mathrm{pert}}$ provides a better approximation of the reproduction number $R$, see SI Fig. LABEL:SI-fig:Rpert_vs_eta-b. This suggests that the advantage of $R_{0}$ for large $\eta$ might be due to a compensatory effect: using $R_{0}$ tends to overestimate the reproduction number, while the network Euler-Lotka equation tends to underestimate $\Lambda$ compared to $\Lambda_{\mathrm{FIT}}$.